# A General and Efficient Federated Split Learning with Pre-trained Image Transformers for Heterogeneous Data

### v1

@@ Abstract
Federated Split Learning (FSL) is a promising distributed learning paradigm in practice, which gathers the strengths of both Federated Learning (FL) and Split Learning (SL) paradigms, to ensure model privacy while diminishing the resource overhead of each client, especially on large transformer models in a resource-constrained environment, e.g., Internet of Things (IoT). However, almost all works merely investigate the performance with simple neural network models in FSL. Despite the minor efforts focusing on incorporating Vision Transformers (ViT) as model architectures, they train ViT from scratch, thereby leading to enormous training overhead in each device with limited resources. Therefore, in this paper, we harness Pre-trained Image Transformers (PITs) as the initial model, coined FES-PIT, to accelerate the training process and improve model robustness. Furthermore, we propose FES-PTZO to hinder the gradient inversion attack, especially having the capability compatible with black-box scenarios, where the gradient information is unavailable. Concretely, FES-PTZO approximates the server gradient by utilizing a zeroth-order (ZO) optimization, which replaces the backward propagation with just one forward process. Empirically, we are the first to provide a systematic evaluation of FSL methods with PITs in real-world datasets, different partial device participations, and heterogeneous data splits. Our experiments verify the effectiveness of our algorithms.
@@ 1 Introduction
  • We incorporate the pre-trained Image Transformer (PIT) into PSL scenarios, called FES-PIT. To further enhance gradient information protection and have the capability compatible with black-box scenarios, we propose FES-PTZO, which leverages zeroth-order (ZO) optimization into FES-PIT.
@@ 3 Methodology
In this section, we first define the problem setup for FSL with pre-trained image Transformers. After that, we elaborate our algorithms: FES-PIT and FES-PTZO.
@@ 3.1 Problem Setup
6                   $h_{c}^{k,t}(i)\leftarrow\textbf{ClientHead}(i,t)$. $\triangleright$ Extract the intermediate feature from local data $b_{s}^{k,t}(i)\leftarrow\mathcal{E}(h_{c}^{k,t}(i))$. $\triangleright$ Generate the smashed feature by using the server body model $\mathcal{L}_{c}^{k,t}(i)\leftarrow\textbf{ClientTail}(i,b_{s}^{k,t}(i))$. $\triangleright$ Finish the forward pass and get the prediction $(w^{i,k+1}_{\mathcal{H}},w^{i,k+1}_{\mathcal{T}})\leftarrow\textbf{ClientUpdate}(i,\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial h_{c}^{k,t}(i)})$. $\triangleright$ Start to backward propagation if $k=K$ then
7                         Option I: (FES-PIT) Generate and store the server gradient by averaging gradients: $g_{s}^{t}(i)=\frac{\partial\mathcal{L}_{c}^{K,t}(i)}{\partial w_{\mathcal{E}}^{t}}$. Option II: (FES-PTZO) Generate and store the server gradient approximation by adopting ZO optimization:
@@
14      Update the weights of the server model for fine-tuning: $w_{\mathcal{E}}^{i,t+1}\leftarrow w_{\mathcal{E}}^{i,t}-\frac{\eta}{m}\sum_{i=1}^{m}g_{s}^{t}(i)$. if t % fedround = 0 then
@@ 2 ClientHead:
4 $h_{c}^{k,t}(i)\leftarrow\mathcal{H}_{i}(\xi_{c}^{k,t}(i))$. Return $h_{c}^{k,t}(i)$. ClientTail:
@@
6 Compute loss value $\mathcal{L}_{c}^{k,t}(i)\leftarrow l_{c}(y_{c}^{k,t}(i),\mathcal{T}_{i}(b_{s}^{k,t}(i)))$. Return $\mathcal{L}_{c}^{k,t}(i)$. ClientUpdate:
$w_{\mathcal{H}}^{i,k+1}\leftarrow w_{\mathcal{H}}^{i,k}-\eta\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{H}}^{i,k}}$. $\triangleright$ Head back-propagation $w_{\mathcal{T}}^{i,k+1}\leftarrow w_{\mathcal{T}}^{i,k}-\eta\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{T}}^{i,k}}$. $\triangleright$ Tail back-propagation Return $(w^{i,k+1}_{\mathcal{H}},w^{i,k+1}_{\mathcal{T}})$.
@@ 3.2 Algorithmic Details
In this subsection, we present our two algorithms: FES-PIT and FES-PTZO. Our goal is to simultaneously optimize a server-side model (an encoder $\mathcal{E}$ of a pre-trained image transformer) and multiple client-side models in terms of head $\mathcal{H}_{i}$ and tail $\mathcal{T}_{i}$ networks to make local clients $i\in\{1,2,...,M\}$ capable of training the large transformer model under the constrained edge resources. Meanwhile, local models suitable for different heterogeneous data can benefit from the pre-trained image transformer, thereby accelerating the training process and reducing the training overhead. Below, we elaborate on the detailed procedure of our algorithms. The paradigm is introduced in Algorithm 1 and 2.
@@
Forward Propagation. During the forward propagation in both FES-PIT and FES-PTZO algorithms, at the beginning of each communication round, each client random initializes the local weights $(w_{\mathcal{H}}^{i},w_{\mathcal{T}}^{i})$ of the head and tail pairs and the encoder weights $w_{\mathcal{E}}$ at the server side retain the weights of the pre-trained image transformer model. Then several clients are sampled to participate in this round and perform their local iteration updates. For each sampled client $i$, it calculates the intermediate feature $h_{c}^{k,t}(i)$ from its local minibatch data by using their local head model $\mathcal{H}_{i}$, which means each client executes the forward propagation. After that, all participating clients transmit their intermediate features to the server side.
@@
To be specific, for FES-PIT, the target is to minimize the objective function (2). So for client $i$, we can generate the gradients of the local head and tail network, which is $\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{H}}^{i,k}}$ and $\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{T}}^{i,k}}$, respectively. After that, the gradients of these local gradients are transmitted to the server. Subsequently, the transformer encoder computes the averaged gradient:
@@
For FES-PTZO, the main focus is on the privacy of the model weights and gradient information in practice, such as black-box scenarios, we adopt the zero-order perturbation following the Simultaneous Perturbation Stochastic Approximation (SPSA) [42, 43] to approximate the high-dimensional gradient efficiently. For gradient approximation, let the perturbation scale is $\epsilon\in(0,1)$ and randomly sampled is $z\in\mathbb{R}^{d}$ with $z\sim\mathcal{N}(0,\mathbf{I}_{d})$. Thus, we obtain the server gradient by:
$g_{s}^{t}(i)\doteq\frac{\mathcal{L}_{c}^{K,t}(i)(w_{\mathcal{E}}^{i,t}+\epsilon z)-\mathcal{L}_{c}^{K,t}(i)(w_{\mathcal{E}}^{i,t}-\epsilon z)}{2\epsilon}z.$ (5)
Note that the primary backward propagation procedure is identical to that in FES-PIT. The sole difference is how to estimate the transformer encoder gradient on the server side due to some important concerns, such as model privacy.
@@ 4.1 Experimental Setup
FES-PIT $96.28_{\pm.13}$ $\textbf{96.87}_{\pm.20}$ $94.77_{\pm.40}$ $\textbf{97.54}_{\pm.25}$ $\textbf{85.55}_{\pm.46}$ $\textbf{81.49}_{\pm.25}$ $93.71_{\pm.75}$ $93.10_{\pm.17}$
FES-PTZO $\textbf{96.93}_{\pm.10}$ $95.88_{\pm.12}$ $\textbf{96.63}_{\pm.30}$ $96.85_{\pm.21}$ $83.05_{\pm.41}$ $80.51_{\pm.13}$ $\textbf{96.77}_{\pm.30}$ $\textbf{94.02}_{\pm.25}$
@@ 4.2 Performance Evaluation
Comparison with the baselines. As shown in Table 2 and Figure 1, the proposed FES-PIT and FES-PTZO outperform other baseline methods with the best stability and better performance in both different datasets and different data heterogeneity scenarios. Specifically, on the CIFAR-10 dataset, FES-PIT and FES-PTZO achieve 96.87% and 95.88% on the Directlet-0.3 setting, 9.19% and 8.20% ahead of the best-compared method Fed-RoD. On the CIFAR-100 dataset, FES-PTZO achieves at least 14.27% and 19.43% improvement from the other baselines on the Pathological-5 and Pathological-10 settings. We attribute that FES-PIT and FES-PTZO have very superior performance and are considerably suitable for the FSL setting based on the performance advantage of PITs. It is also observed that FES-PTZO has better performance while protecting the model privacy on the server side, demonstrating that ZO optimization is capable of applying in distributed learning, such as FSL.
Evaluation on various non-IID scenarios. We discuss two types of data heterogeneity: Dirichlet distribution and Pathological distribution under various pre-trained transformer models on the CIFAR-10 dataset in Table 3, and prove the effectiveness and robustness of the proposed methods. In Dirichlet distribution, since the local training can’t cater for all classes inside clients, the accuracy decreases with the level of heterogeneity decreasing. It is clearly seen that our algorithms: FES-PIT and FES-PTZO can achieve stronger stability for several heterogeneous settings under various pre-trained transformer models. Pathological distribution defines limited classes for each client which is a higher level of heterogeneity. FES-PIT and FES-PTZO still have better performance on CIFAR-10 with only 2 categories per client.
@@ 5 Conclusion
Supplementary Material for “ A General and Efficient Federated Split Learning with Pre-trained Image Transformers for Heterogeneous Data ”
@@ 0.A.1 Datasets and Data Partition
0.A.2 More details about the Fed-PIT and Fed-PTZO
For Fed-PIT and Fed-PTZO, we train the model for 5 epochs per round as the same as other baselines. Therefore, we conduct ablation experiments to select the most appropriate parameters for each data distribution. We set fedround to 50. Model parameters are optimized by Adam with a learning rate decay of 0.0002. The weight perturbation ratio in Fed-PTZO is set to zero in the experiments. Furthermore, in Fed-PTZO, the main hyper-parameters of ZO optimization are set to $\epsilon=0.0001$, the learning rate of the server-side model $lr_{server}=1e-6$ with the decaying rate 0.5 and the learning rate of the client model $lr_{client}=2e-4$.
@@ 0.A.3 More Details about Baselines
Fed-PIT 89.35 $\pm$ 0.30 87.77 $\pm$ 0.08 96.77 $\pm$ 0.10 95.70 $\pm$ 0.50
Fed-PTZO 87.10 $\pm$ 0.11 85.40 $\pm$ 0.50 96.83 $\pm$ 0.16 96.41 $\pm$ 0.33
Comparison with the baselines. As shown in Table 5 and Figure 4, the proposed FES-PIT and FES-PTZO outperform other baseline methods with the best stability and better performance in both different datasets and different data heterogeneity scenarios. Specifically, FES-PIT and FES-PTZO achieve 87.77% and 85.40% on the Directlet-0.3 setting, which are very far ahead of the best-compared method Fed-RoD. We attribute that FES-PIT and FES-PTZO have very superior performance and are considerably suitable for the FSL setting based on the performance advantage of PITs.
Evaluation on various non-IID scenarios. We discuss two types of data heterogeneity: Dirichlet distribution and Pathological distribution under various pre-trained transformer models in Table 5, and prove the effectiveness and robustness of the proposed methods. In Dirichlet distribution, since the local training can’t cater for all classes inside clients, the accuracy decreases with the level of heterogeneity decreasing. It is clearly seen that our algorithms: FES-PIT and FES-PTZO can achieve stronger stability for several heterogeneous settings under various pre-trained transformer models.

### v2

# Heterogeneous Federated Learning with Splited Language Model

@@ Abstract
Federated Split Learning (FSL) is a promising distributed learning paradigm in practice, which gathers the strengths of both Federated Learning (FL) and Split Learning (SL) paradigms, to ensure model privacy while diminishing the resource overhead of each client, especially on large transformer models in a resource-constrained environment, e.g., Internet of Things (IoT). However, almost all works merely investigate the performance with simple neural network models in FSL. Despite the minor efforts focusing on incorporating Vision Transformers (ViT) as model architectures, they train ViT from scratch, thereby leading to enormous training overhead in each device with limited resources. Therefore, in this paper, we harness Pre-trained Image Transformers (PITs) as the initial model, coined FedV, to accelerate the training process and improve model robustness. Furthermore, we propose FedVZ to hinder the gradient inversion attack, especially having the capability compatible with black-box scenarios, where the gradient information is unavailable. Concretely, FedVZ approximates the server gradient by utilizing a zeroth-order (ZO) optimization, which replaces the backward propagation with just one forward process. Empirically, we are the first to provide a systematic evaluation of FSL methods with PITs in real-world datasets, different partial device participations, and heterogeneous data splits. Our experiments verify the effectiveness of our algorithms.
@@ 1 Introduction
  • We incorporate the pre-trained Image Transformer (PIT) into PSL scenarios, called FedV. To further enhance gradient information protection and have the capability compatible with black-box scenarios, we propose FedVZ, which leverages zeroth-order (ZO) optimization into FedV.
@@ 3 Methodology
In this section, we first define the problem setup for FSL with pre-trained image Transformers. After that, we elaborate our algorithms: FedV and FedVZ.
@@ 3.1 Problem Setup
6                   $h_{c}^{k,t}(i)\leftarrow\textbf{ClientHead}(i,t)$. $\triangleright$ Extract the intermediate feature from local data
7                  $b_{s}^{k,t}(i)\leftarrow\mathcal{E}(h_{c}^{k,t}(i))$. $\triangleright$ Generate the smashed feature by using the server body model
8                  $\mathcal{L}_{c}^{k,t}(i)\leftarrow\textbf{ClientTail}(i,b_{s}^{k,t}(i))$. $\triangleright$ Finish the forward pass and get the prediction
9                  $(w^{i,k+1}_{\mathcal{H}},w^{i,k+1}_{\mathcal{T}})\leftarrow\textbf{ClientUpdate}(i,\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial h_{c}^{k,t}(i)})$. $\triangleright$ Start to backward propagation
10                  if $k=K$ then
11                         Option I: (FedV) Generate and store the server gradient by averaging gradients: $g_{s}^{t}(i)=\frac{\partial\mathcal{L}_{c}^{K,t}(i)}{\partial w_{\mathcal{E}}^{t}}$.
12                        Option II: ( FedVZ) Generate and store the server gradient approximation by adopting ZO optimization:
@@
19      Update the weights of the server model for fine-tuning: $w_{\mathcal{E}}^{i,t+1}\leftarrow w_{\mathcal{E}}^{i,t}-\frac{\eta}{m}\sum_{i=1}^{m}g_{s}^{t}(i)$.
20      if t % fedround = 0 then
@@ 2 ClientHead:
4 $h_{c}^{k,t}(i)\leftarrow\mathcal{H}_{i}(\xi_{c}^{k,t}(i))$.
5Return $h_{c}^{k,t}(i)$.
6ClientTail:
@@
8 Compute loss value $\mathcal{L}_{c}^{k,t}(i)\leftarrow l_{c}(y_{c}^{k,t}(i),\mathcal{T}_{i}(b_{s}^{k,t}(i)))$.
9Return $\mathcal{L}_{c}^{k,t}(i)$.
10ClientUpdate:
11 $w_{\mathcal{H}}^{i,k+1}\leftarrow w_{\mathcal{H}}^{i,k}-\eta\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{H}}^{i,k}}$. $\triangleright$ Head back-propagation
[…]
Return $(w^{i,k+1}_{\mathcal{H}},w^{i,k+1}_{\mathcal{T}})$.
@@ 3.2 Algorithmic Details
In this subsection, we present our two algorithms: FedV and FedVZ. Our goal is to simultaneously optimize a server-side model (an encoder $\mathcal{E}$ of a pre-trained image transformer) and multiple client-side models in terms of head $\mathcal{H}_{i}$ and tail $\mathcal{T}_{i}$ networks to make local clients $i\in\{1,2,...,M\}$ capable of training the large transformer model under the constrained edge resources. Meanwhile, local models suitable for different heterogeneous data can benefit from the pre-trained image transformer, thereby accelerating the training process and reducing the training overhead. Below, we elaborate on the detailed procedure of our algorithms. The paradigm is introduced in Algorithm 1 and 2.
@@
Forward Propagation. During the forward propagation in both FedV and FedVZ algorithms, at the beginning of each communication round, each client random initializes the local weights $(w_{\mathcal{H}}^{i},w_{\mathcal{T}}^{i})$ of the head and tail pairs and the encoder weights $w_{\mathcal{E}}$ at the server side retain the weights of the pre-trained image transformer model. Then several clients are sampled to participate in this round and perform their local iteration updates. For each sampled client $i$, it calculates the intermediate feature $h_{c}^{k,t}(i)$ from its local minibatch data by using their local head model $\mathcal{H}_{i}$, which means each client executes the forward propagation. After that, all participating clients transmit their intermediate features to the server side.
@@
To be specific, for FedV, the target is to minimize the objective function (2). So for client $i$, we can generate the gradients of the local head and tail network, which is $\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{H}}^{i,k}}$ and $\frac{\partial\mathcal{L}_{c}^{k,t}(i)}{\partial w_{\mathcal{T}}^{i,k}}$, respectively. After that, the gradients of these local gradients are transmitted to the server. Subsequently, the transformer encoder computes the averaged gradient:
@@
For FedVZ, the main focus is on the privacy of the model weights and gradient information in practice, such as black-box scenarios, we adopt the zero-order perturbation following the Simultaneous Perturbation Stochastic Approximation (SPSA) [42, 43] to approximate the high-dimensional gradient efficiently. For gradient approximation, let the perturbation scale is $\epsilon\in(0,1)$ and randomly sampled is $z\in\mathbb{R}^{d}$ with $z\sim\mathcal{N}(0,\mathbf{I}_{d})$. Thus, we obtain the server gradient by:
$g_{s}^{t}(i)\doteq\frac{\mathcal{L}_{c}^{K,t}(w_{\mathcal{E}}^{i,t}+\epsilon z)-\mathcal{L}_{c}^{K,t}(w_{\mathcal{E}}^{i,t}-\epsilon z)}{2\epsilon}z.$ (5)
Note that the primary backward propagation procedure is identical to that in FedV. The sole difference is how to estimate the transformer encoder gradient on the server side due to some important concerns, such as model privacy.
@@ 4.1 Experimental Setup
FedV $96.28_{\pm.13}$ $\textbf{96.87}_{\pm.20}$ $94.77_{\pm.40}$ $\textbf{97.54}_{\pm.25}$ $\textbf{85.55}_{\pm.46}$ $\textbf{81.49}_{\pm.25}$ $93.71_{\pm.75}$ $93.10_{\pm.17}$
FedVZ $\textbf{96.93}_{\pm.10}$ $95.88_{\pm.12}$ $\textbf{96.63}_{\pm.30}$ $96.85_{\pm.21}$ $83.05_{\pm.41}$ $80.51_{\pm.13}$ $\textbf{96.77}_{\pm.30}$ $\textbf{94.02}_{\pm.25}$
@@ 4.2 Performance Evaluation
Comparison with the baselines. As shown in Table 2 and Figure 1, the proposed FedV and FedVZ outperform other baseline methods with the best stability and better performance in both different datasets and different data heterogeneity scenarios. Specifically, on the CIFAR-10 dataset, FedV and FedVZ achieve 96.87% and 95.88% on the Directlet-0.3 setting, 9.19% and 8.20% ahead of the best-compared method Fed-RoD. On the CIFAR-100 dataset, FedVZ achieves at least 14.27% and 19.43% improvement from the other baselines on the Pathological-5 and Pathological-10 settings. We attribute that FedV and FedVZ have very superior performance and are considerably suitable for the FSL setting based on the performance advantage of PITs. It is also observed that FedVZ has better performance while protecting the model privacy on the server side, demonstrating that ZO optimization is capable of applying in distributed learning, such as FSL.
Evaluation on various non-IID scenarios. We discuss two types of data heterogeneity: Dirichlet distribution and Pathological distribution under various pre-trained transformer models on the CIFAR-10 dataset in Table 3, and prove the effectiveness and robustness of the proposed methods. In Dirichlet distribution, since the local training can’t cater for all classes inside clients, the accuracy decreases with the level of heterogeneity decreasing. It is clearly seen that our algorithms: FedV and FedVZ can achieve stronger stability for several heterogeneous settings under various pre-trained transformer models. Pathological distribution defines limited classes for each client which is a higher level of heterogeneity. FedV and FedVZ still have better performance on CIFAR-10 with only 2 categories per client.
@@ 5 Conclusion
Supplementary Material for “Heterogeneous Federated Learning with Splited Language Model”
@@ 0.A.1 Datasets and Data Partition
0.A.2 More details about the FedV and FedVZ
For FedV and FedVZ, we train the model for 5 epochs per round as the same as other baselines. Therefore, we conduct ablation experiments to select the most appropriate parameters for each data distribution. We set fedround to 50. Model parameters are optimized by Adam with a learning rate decay of 0.0002. The weight perturbation ratio in FedVZ is set to zero in the experiments. Furthermore, in FedVZ, the main hyper-parameters of ZO optimization are set to $\epsilon=0.0001$, the learning rate of the server-side model $lr_{server}=1e-6$ with the decaying rate 0.5 and the learning rate of the client model $lr_{client}=2e-4$.
@@ 0.A.3 More Details about Baselines
FedV 89.35 $\pm$ 0.30 87.77 $\pm$ 0.08 96.77 $\pm$ 0.10 95.70 $\pm$ 0.50
FedVZ 87.10 $\pm$ 0.11 85.40 $\pm$ 0.50 96.83 $\pm$ 0.16 96.41 $\pm$ 0.33
Comparison with the baselines. As shown in Table 5 and Figure 4, the proposed FedV and FedVZ outperform other baseline methods with the best stability and better performance in both different datasets and different data heterogeneity scenarios. Specifically, FedV and FedVZ achieve 87.77% and 85.40% on the Directlet-0.3 setting, which are very far ahead of the best-compared method Fed-RoD. We attribute that FedV and FedVZ have very superior performance and are considerably suitable for the FSL setting based on the performance advantage of PITs.
Evaluation on various non-IID scenarios. We discuss two types of data heterogeneity: Dirichlet distribution and Pathological distribution under various pre-trained transformer models in Table 5, and prove the effectiveness and robustness of the proposed methods. In Dirichlet distribution, since the local training can’t cater for all classes inside clients, the accuracy decreases with the level of heterogeneity decreasing. It is clearly seen that our algorithms: FedV and FedVZ can achieve stronger stability for several heterogeneous settings under various pre-trained transformer models.